Stoney Brooke, Inc., has sales of $1,120,000 and cost of goods sold of $1,017,900. The firm had a beginning inventory of $51,000 and an ending inventory of $66,000. What is the length of the inventory period? Assume 365 days per year.
Answer #1

Given,

Cost of goods sold = $1017900

Beginning inventory = $51000

Ending inventory = $66000

Average period inventory = (beg+end)/2 = (51000+66000)/2 = $58500

Inventory turnover ratio = COGS/Avg. inventory = 1017900/58500 = 17.4

Avg. Inventory period = Number of days in period/Inventory turnover ratio = 365/17.4 = 20.98 days

So, length of the inventory period is 20.98 days